# Functoriality in Finitary Vacuum Einstein Gravity and Free Yang-Mills

### 24 Jan 2024 v1

@@ 1.1.1 Sorkin’s Finitary Posets
The collection $\overleftarrow{\mathcal{P}}=(\mathcal{P}_{i})_{i\in I}$ of such finposets is seen to constitute a so-called inverse or projective system, or net, of posets, which is seen to have a projective limit space effectively homeomorphic to the continuous $\mathcal{C}^{0}$-manifold $M$ [^2] The index $i\in I$ in $\mathcal{P}_{i}$ is the so-called refinement net index, whereby $\mathcal{U}_{i}\prec\mathcal{U}{j}$ (reads: the open cover $\mathcal{U}{i}$ is finer than the open cover $\mathcal{U}{j}$, and conversely, $\mathcal{U}{j}$ is coarser than $\mathcal{U}{i}$) if $\mathcal{U}_{i}$ has ‘smaller’ and more numerous open sets than $\mathcal{U}_{j}$. The aforementioned inverse limit now reads: the $\mathcal{C}^{0}$-manifold is recovered at the inverse limit of infinite refinement of the open covers $\mathcal{U}_{i}$s in the net (as $i\rightarrow\infty$).
@@ 1.2.1 Briefly revisiting $\mathbf{A}$-connections in ADG
noindent which is also defined functorially as an $\mathbf{A}$-linear, Leibnizian sheaf morphism, acting on the relevant module sheaves as follows:
@@ 1.2.2 First encounter with ‘Dynamical’ Functoriality: ADG-theoretic Fields
  • From a geometric (pre)quantization perspective, the local sections of $\mathbf{\mathcal{E}}$ correspond to quantum particle states [40, 45, 44, 51, 54, 55, 72, 73]. If $\mathbf{\mathcal{E}}$ is a line sheaf $\mathcal{L}$ (:a vector sheaf of rank $n=1$), its local sections represent the quantum particle states of a boson like the ‘photon’, hence the ADG-field $\mathcal{F}_{Max}=(\mathcal{D},\mathcal{L})$ is coined the Maxwell field. More general Yang-Mills ADG-fields are represented as connections $\mathcal{D}$ on vector sheaves $\mathbf{\mathcal{E}}$ of rank $n>1$. ADG-theoretically, we represent them by the pair: $\mathcal{F}_{YM}=(\mathcal{D},\mathbf{\mathcal{E}})$. Finally, the gravitational connections constitute ADG-theoretic Einstein fields: $\mathcal{F}_{Einst}=(\mathcal{D},\mathbf{\mathcal{E}})$.
@@
  • The Maxwell $\mathcal{F}_{Max}=(\mathcal{D},\mathcal{L})$, Yang-Mills $\mathcal{F}_{YM}=(\mathcal{D},\mathbf{\mathcal{E}})$ and Einstein $\mathcal{F}_{Einst}=(\mathcal{D},\mathbf{\mathcal{E}})$ ADG-fields can be organised, as objects, into respective categories with categorical sheaf morphisms as arrows between them, coined: the Maxwell Category $\mathcal{T}_{Max}$, the Yang-Mills Category $\mathcal{T}_{YM}$ and the Einstein Category $\mathcal{T}_{Einst}$ categories [44, 49, 51]. In contradistinction to the ‘flat’, ungauged and ‘static-kinematical’ categories of differential triads that we alluded to earlier [64, 65, 44], these three categories are ‘dynamical’ in character, in the sense that object-fields in them obey and satisfy certain dynamical laws of motion [^20] The Maxwell, Yang-Mills and Einstein dynamical (differential) equations, which in turn derive from Lagrangian variation of corresponding action functionals. Read on. and the $\mathbf{A}$-functoriality thereof corresponds to the $\mathbf{A}$-invariance and the ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ local gauge/generalised coordinates’ invariance of the respective dynamical laws of motion [49, 51]. [^21] We shall return to discuss further the categorical implications and the deeper physical interpretation of the generalised ADG-theoretic conception of local gauge invariance as $\mathbf{A}$-invariance and the $\otimes_{\mathbf{A}}$-functoriality of the ADG-theoretic gauge dynamics in the sequel.
@@
The discussion above brings us to the all-important issue of gauge covariance, local gauge invariance and their intimate relation to the basic ADG-theoretic notion of $\mathcal{A}$-invariance. [^23] As it has also been observed in past publications [53, 54, 55], in our work we use the symbol $\mathbf{A}$ for the structure sheaf of algebras of generalised ADG-theoretic coordinate functions, as opposed to $\mathcal{A}$ used throughout Mallios’s work [38, 39, 44, 49, 51], as we have reserved the symbol $\mathcal{A}$ for local Einstein gravitational and Yang-Mills local gauge potentials as in (13). In this subsection, we will focus only on [non-abelian] Einstein and Yang-Mills ADG-fields on higher rank vector sheaves, leaving the abelian case of Maxwell fields on line sheaves to their exhaustive treatment in the monograph references [38, 44]. [^24] The epithets abelian and non-abelian above pertain, as in the usual theory [23], to the structure gauge groups being commutative and non-commutative, respectively. Indeed, the principal group sheaves associated with the line sheaves $\mathcal{L}$ of the ADG-theoretic Maxwell fields $\mathcal{F}_{Max}$ above carry abelian unitary (:$\equiv U_{1}(\mathbf{A})$) groups in their stalks, while as we saw couple of footnotes earlier, the structure group sheaf ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ of dynamical automorphisms of the vector sheaf $\mathbf{\mathcal{E}}$ of, say, the Einstein field $\mathcal{F}_{Einst}$ is locally homomorphic to $M_{n}(\mathbf{A}(U))$, which is manifestly noncommutative.
@@ 1.2.5 Local Gauge Transformation of $\mathcal{D}$
which, plainly, is a local (i.e., relative to $U\cap V$) section of the ‘natural’ structure group sheaf $\mathcal{G}\mathcal{L}(n,\mathbf{A})$ of $\mathbf{\mathcal{E}}$ [^32] As noted earlier, one may recognise $\mathcal{G}\mathcal{L}(n,\mathbf{A})$ above as the local version of the automorphism principal group sheaf ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ of $\mathbf{\mathcal{E}}$. The adjective ‘local’ here pertains to the fact mentioned earlier that ADG assumes that $\mathbf{\mathcal{E}}$ is locally isomorphic to $\mathbf{A}^{n}$.—that is to say, $g_{ij}\in\mathrm{GL}(n,\mathbf{A}(U\cap V))=\mathcal{G}\mathcal{L}(n,\mathbf{A})(U\cap V)$.
@@ 1.3.1 ADG-Kinematics: The Affine Space of $\mathbf{A}$-Connections
As also highlighted in [55], it follows that the generalised ‘kinematical’ space of the theory is the affine space $\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})$ of $\mathbf{A}$-connections $\mathcal{A}$ on $\mathbf{\mathcal{E}}$. Moreover, since the Lagrangians involved in (25) are invariant under the group sheaf $\mathcal{G}(\mathbf{\mathcal{E}})={\mathcal{A}}ut\mathbf{\mathcal{E}}$ of local automorphisms (:local gauge transformations) of $\mathbf{\mathcal{E}}$,
the relevant kinematical space is the moduli space $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ of gauge equivalent $\mathbf{A}$-connections $\mathcal{A}$ on $\mathbf{\mathcal{E}}$ [^36] Recall from [55] that $\mathcal{G}(\mathbf{\mathcal{E}})|_{U}={\mathcal{A}}ut\mathbf{\mathcal{E}}|_{U}=(\mathrm{Aut}_{\mathbf{A}}\mathbf{\mathcal{E}})(U):=\Gamma(U,{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}})\equiv M^{n}_{U}(\mathbf{A})$. Thus, $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ is the so-called $\mathcal{G}$-orbit space as the structure gauge group sheaf ${\mathcal{A}}ut\mathbf{\mathcal{E}}(\mathbf{A})$ cuts through the affine space $\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})$, carving out ‘paths’ or ‘orbits’ of gauge equivalent connections in the process, which, in turn, leave the corresponding curvature Lagrangians in (25) invariant under (local) gauge transformations.
Hence the integration sign in the dynamical action functionals in (25), which supposedly extends over the base topological space $X$, in effect extends over the moduli space $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ of gauge equivalent connections. [^37] From [55, 49, 51, 44] we read that a suitably defined ADG-theoretic Radon-type of $\mathbf{A}$-linear continuous integration measure $d\mu$ on a suitably topologised $\textsf{A}/\mathcal{G}$ ($d\mu:\;\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}\longrightarrow\mathbf{A}$) is expected to render rigorous the dynamical action integrals in (25).
@@ 1.3.2 $\mathbf{A}$-Invariance and ‘Dynamical’ Functoriality
“…Therefore, one thus realizes that, the fundamental adjunction [^46] Our geometric morphism $\mathcal{G}\mathcal{M}$ in equations (27) and (28) earlier., preserves the $\mathbf{A}$-invariance, for any $\mathbf{A}$-invariant function, [^47] Especially, for the $\mathbf{A}$-invariant Einstein and Yang-Mills action functionals in (25) above, which, as we saw earlier, are $\mathbf{A}$-valued functionals defined on the moduli space $\textsf{A}_{\mathbf{A}}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut\mathbf{\mathcal{E}}$ of gauge equivalent $\mathbf{A}$-connections on $\mathbf{\mathcal{E}}$. pertaining to the two basic functors appearing in the aforesaid adjunction…” [^48] That is, the homological $\mathrm{Hom}_{\mathbf{A}}$ and $\otimes_{\mathbf{A}}$ adjoint functors constituting $\mathcal{G}\mathcal{M}_{\mathbf{A}}$ in equations (27) and (28) earlier.
@@
Corollary to the Fundamental Theorem of ADG-Gauge Theory. The algebra (sheaf) of our generalised coordinate measurements $\mathbf{A}$ respects (i.e. it is ‘non-perturbing’ and it leaves invariant) the functorial ADG-theoretic gauge field dynamics for Vacuum Einstein Gravity and Free Yang-Mills theories, hence, in return, it entails and almost ‘mandates’ that the (local) Relativity Group (sheaf) of the theory is $\mathcal{G}\equiv{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ [^51] That the local relativity group in ADG-GT is ‘naturally’ ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ (:see subsection next) has been amply expounded in [53, 54, 55, 44, 49, 50, 69, 70, 71, 72, 73]. Thus, $\mathbf{A}$-invariance, via the ‘dynamical’ geometric morphism $\mathcal{G}\mathcal{M}_{\mathbf{A}}$, which in turn entails the $\mathcal{G}={\mathcal{A}}ut\mathbf{\mathcal{E}}$-gauge invariance of the ADG-field dynamics (expressed via the geometric morph of the connection field—the field’s curvature), corresponds to a dynamical version of the Kleinian conception of geometry. [^52] According to Felix Klein, ‘the geometry of an object is all the transformations of it that leave it invariant[32].
@@ 2.1.1 The Functoriality of 3rd Quantization in ADG-GT
$[\mathbf{\mathcal{E}},\mathcal{D}]={\mathcal{A}}ut\mathbf{\mathcal{E}}\,\stackrel{{\scriptstyle\mathrm{locally}}}{{\longrightarrow}}\,[s\in\mathbf{\mathcal{E}}(U),d+\mathcal{A}]=M_{n}(\mathbf{A}_{U})s=M_{n}(\mathbf{A}(U))s$ (37)
@@
The epithet ‘unitary’ for $\mathbf{U}$ in (39) above indicates that it is a holistic entity, an inseparable whole, encoding state ($\mathbf{\mathcal{E}}$), dynamical changes of state ($\mathcal{D}$), gauge symmetries and invariances of dynamical changes of state (${\mathcal{A}}ut\mathbf{\mathcal{E}}$), and quantum uncertainty of ‘determination’ ($\mathcal{Q}$), all-4-in-1.

### 27 Jan 2024 v2

@@ 1.1.1 Sorkin’s Finitary Posets
The collection $\overleftarrow{\mathcal{P}}=(\mathcal{P}_{i})_{i\in I}$ of such finposets is seen to constitute a so-called inverse or projective system, or net, of posets, which is seen to have a projective limit space effectively homeomorphic to the continuous $\mathcal{C}^{0}$-manifold $M$ [^2] The index $i\in I$ in $\mathcal{P}_{i}$ is the so-called refinement net index, whereby $\mathcal{U}_{i}\prec\mathcal{U}{j}$ (reads: the open cover $\mathcal{U}_{j}$ is finer than the open cover $\mathcal{U}_{i}$, and conversely, $\mathcal{U}_{i}$ is coarser than $\mathcal{U}_{j}$) if $\mathcal{U}_{j}$ has ‘smaller’ and more numerous open sets than $\mathcal{U}_{i}$. The aforementioned inverse limit now reads: the $\mathcal{C}^{0}$-manifold is recovered at the inverse limit of infinite refinement of the open covers $\mathcal{U}_{i}$s in the net (as $i\rightarrow\infty$).
@@ 1.2.1 Briefly revisiting $\mathbf{A}$-connections in ADG
which is also defined functorially as an $\mathbf{A}$-linear, Leibnizian sheaf morphism, acting on the relevant module sheaves as follows:
@@ 1.2.2 First encounter with ‘Dynamical’ Functoriality: ADG-theoretic Fields
  • From a geometric (pre)quantization perspective, the local sections of $\mathbf{\mathcal{E}}$ correspond to quantum particle states [40, 45, 44, 51, 54, 55, 72, 73]. If $\mathbf{\mathcal{E}}$ is a line sheaf $\mathcal{L}$ (:a vector sheaf of rank $n=1$), its local sections represent the quantum particle states of a boson like the ‘photon’, hence the ADG-field $\mathcal{F}_{Max}=(\mathcal{L},\mathcal{D})$ is coined the Maxwell field. More general Yang-Mills ADG-fields are represented as connections $\mathcal{D}$ on vector sheaves $\mathbf{\mathcal{E}}$ of rank $n>1$. ADG-theoretically, we represent them by the pair: $\mathcal{F}_{YM}=(\mathbf{\mathcal{E}},\mathcal{D})$. Finally, the gravitational connections constitute ADG-theoretic Einstein fields: $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$.
@@
  • The Maxwell $\mathcal{F}_{Max}=(\mathcal{L},\mathcal{D})$, Yang-Mills $\mathcal{F}_{YM}=(\mathbf{\mathcal{E}},\mathcal{D})$ and Einstein $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ ADG-fields can be organised, as objects, into respective categories with categorical sheaf morphisms as arrows between them, coined: the Maxwell Category $\mathcal{T}_{Max}$, the Yang-Mills Category $\mathcal{T}_{YM}$ and the Einstein Category $\mathcal{T}_{Einst}$ categories [44, 49, 51]. In contradistinction to the ‘flat’, ungauged and ‘static-kinematical’ categories of differential triads that we alluded to earlier [64, 65, 44], these three categories are ‘dynamical’ in character, in the sense that object-fields in them obey and satisfy certain dynamical laws of motion [^20] The Maxwell, Yang-Mills and Einstein dynamical (differential) equations, which in turn derive from Lagrangian variation of corresponding action functionals. Read on. and the $\mathbf{A}$-functoriality thereof corresponds to the $\mathbf{A}$-invariance and the ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ local gauge/generalised coordinates’ invariance of the respective dynamical laws of motion [49, 51]. [^21] We shall return to discuss further the categorical implications and the deeper physical interpretation of the generalised ADG-theoretic conception of local gauge invariance as $\mathbf{A}$-invariance and the $\otimes_{\mathbf{A}}$-functoriality of the ADG-theoretic gauge dynamics in the sequel.
@@
The discussion above brings us to the all-important issue of gauge covariance, local gauge invariance and their intimate relation to the basic ADG-theoretic notion of $\mathcal{A}$-invariance. [^23] As it has also been observed in past publications [53, 54, 55], in our work we use the symbol $\mathbf{A}$ for the structure sheaf of algebras of generalised ADG-theoretic coordinate functions, as opposed to $\mathcal{A}$ used throughout Mallios’s work [38, 39, 44, 49, 51], as we have reserved the symbol $\mathcal{A}$ for local Einstein gravitational and Yang-Mills local gauge potentials as in (13). In this subsection, we will focus only on [non-abelian] Einstein and Yang-Mills ADG-fields on higher rank vector sheaves, leaving the abelian case of Maxwell fields on line sheaves to their exhaustive treatment in the monograph references [38, 44]. [^24] The epithets abelian and non-abelian above pertain, as in the usual theory [23], to the structure gauge groups being commutative and non-commutative, respectively. Indeed, the principal group sheaves associated with the line sheaves $\mathcal{L}$ of the ADG-theoretic Maxwell fields $\mathcal{F}_{Max}$ above carry abelian unitary (:$\equiv U_{1}(\mathbf{A})$) groups in their stalks, while as we saw couple of footnotes earlier, the structure group sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of dynamical automorphisms of the vector sheaf $\mathbf{\mathcal{E}}$ of, say, the Einstein field $\mathcal{F}_{Einst}$ is locally homomorphic to $M_{n}(\mathbf{A}(U))$, which is manifestly noncommutative.
@@ 1.2.5 Local Gauge Transformation of $\mathcal{D}$
which, plainly, is a local (i.e., relative to $U\cap V$) section of the ‘natural’ structure group sheaf $\mathcal{G}\mathcal{L}(n,\mathbf{A})$ of $\mathbf{\mathcal{E}}$ [^32] As noted earlier, one may recognise $\mathcal{G}\mathcal{L}(n,\mathbf{A})$ above as the local version of the automorphism principal group sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of $\mathbf{\mathcal{E}}$. The adjective ‘local’ here pertains to the fact mentioned earlier that ADG assumes that $\mathbf{\mathcal{E}}$ is locally isomorphic to $\mathbf{A}^{n}$.—that is to say, $g_{ij}\in\mathrm{GL}(n,\mathbf{A}(U\cap V))=\mathcal{G}\mathcal{L}(n,\mathbf{A})(U\cap V)$.
@@ 1.3.1 ADG-Kinematics: The Affine Space of $\mathbf{A}$-Connections
As also highlighted in [55], it follows that the generalised ‘kinematical’ space of the theory is the affine space $\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})$ of $\mathbf{A}$-connections $\mathcal{A}$ on $\mathbf{\mathcal{E}}$. Moreover, since the Lagrangians involved in (25) are invariant under the group sheaf $\mathcal{G}(\mathbf{\mathcal{E}})={\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of local automorphisms (:local gauge transformations) of $\mathbf{\mathcal{E}}$,
the relevant kinematical space is the moduli space $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of gauge equivalent $\mathbf{A}$-connections $\mathcal{A}$ on $\mathbf{\mathcal{E}}$ [^36] Recall from [55] that $\mathcal{G}(\mathbf{\mathcal{E}})|_{U}={\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}|_{U}=(\mathrm{Aut}_{\mathbf{A}}\mathbf{\mathcal{E}})(U):=\Gamma(U,{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}})\equiv M^{n}_{U}(\mathbf{A})$. Thus, $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ is the so-called $\mathcal{G}$-orbit space as the structure gauge group sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}(\mathbf{A})$ cuts through the affine space $\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})$, carving out ‘paths’ or ‘orbits’ of gauge equivalent connections in the process, which, in turn, leave the corresponding curvature Lagrangians in (25) invariant under (local) gauge transformations.
Hence the integration sign in the dynamical action functionals in (25), which supposedly extends over the base topological space $X$, in effect extends over the moduli space $\textsf{A}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of gauge equivalent connections. [^37] From [55, 49, 51, 44] we read that a suitably defined ADG-theoretic Radon-type of $\mathbf{A}$-linear continuous integration measure $d\mu$ on a suitably topologised $\textsf{A}/\mathcal{G}$ ($d\mu:\;\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}\longrightarrow\mathbf{A}$) is expected to render rigorous the dynamical action integrals in (25).
@@ 1.3.2 $\mathbf{A}$-Invariance and ‘Dynamical’ Functoriality
“…Therefore, one thus realizes that, the fundamental adjunction [^46] Our geometric morphism $\mathcal{G}\mathcal{M}$ in equations (27) and (28) earlier., preserves the $\mathbf{A}$-invariance, for any $\mathbf{A}$-invariant function, [^47] Especially, for the $\mathbf{A}$-invariant Einstein and Yang-Mills action functionals in (25) above, which, as we saw earlier, are $\mathbf{A}$-valued functionals defined on the moduli space $\textsf{A}_{\mathbf{A}}/\mathcal{G}=\textsf{A}_{\mathbf{A}}(\mathbf{\mathcal{E}})/{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of gauge equivalent $\mathbf{A}$-connections on $\mathbf{\mathcal{E}}$. pertaining to the two basic functors appearing in the aforesaid adjunction…” [^48] That is, the homological $\mathrm{Hom}_{\mathbf{A}}$ and $\otimes_{\mathbf{A}}$ adjoint functors constituting $\mathcal{G}\mathcal{M}_{\mathbf{A}}$ in equations (27) and (28) earlier.
@@
Corollary to the Fundamental Theorem of ADG-Gauge Theory. The algebra (sheaf) of our generalised coordinate measurements $\mathbf{A}$ respects (i.e. it is ‘non-perturbing’ and it leaves invariant) the functorial ADG-theoretic gauge field dynamics for Vacuum Einstein Gravity and Free Yang-Mills theories, hence, in return, it entails and almost ‘mandates’ that the (local) Relativity Group (sheaf) of the theory is $\mathcal{G}\equiv{\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ [^51] That the local relativity group in ADG-GT is ‘naturally’ ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ (:see subsection next) has been amply expounded in [53, 54, 55, 44, 49, 50, 69, 70, 71, 72, 73]. Thus, $\mathbf{A}$-invariance, via the ‘dynamical’ geometric morphism $\mathcal{G}\mathcal{M}_{\mathbf{A}}$, which in turn entails the $\mathcal{G}={\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$-gauge invariance of the ADG-field dynamics (expressed via the geometric morph of the connection field—the field’s curvature), corresponds to a dynamical version of the Kleinian conception of geometry. [^52] According to Felix Klein, ‘the geometry of an object is all the transformations of it that leave it invariant[32].
@@ 2.1.1 The Functoriality of 3rd Quantization in ADG-GT
$[\mathbf{\mathcal{E}},\mathcal{D}]={\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}\,\stackrel{{\scriptstyle\mathrm{locally}}}{{\longrightarrow}}\,[s\in\mathbf{\mathcal{E}}(U),d+\mathcal{A}]=M_{n}(\mathbf{A}_{U})s=M_{n}(\mathbf{A}(U))s$ (37)
@@
The epithet ‘unitary’ for $\mathbf{U}$ in (39) above indicates that it is a holistic entity, an inseparable whole, encoding state ($\mathbf{\mathcal{E}}$), dynamical changes of state ($\mathcal{D}$), gauge symmetries and invariances of dynamical changes of state (${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$), and quantum uncertainty of ‘determination’ ($\mathcal{Q}$), all-4-in-1.